sandra baker held a $3500, 70-day note dated March 14, bearing interest at 10%. on april 14, she took the note to Loyal Capital Bank which discounted it at 15%. Use a 365 day year to find the interest amount of the loan
Answer #1

Solution:

Period of maturity remaining on April 14 for note = 70 - 31 = 39 days

Interest amount on the loan = $3,500*15%*39/365 = $56.10